Solve the equation sin? + 2cos? - 2 = 0 for 0 = ? = 2π.
Answer by Alan3354(21605)

Solve the equation sin? + 2cos? - 2 = 0 for 0 = ? = 2π.
--------------
sin = 2 - 2cos
sqrt(1-cos^2) = 2 - 2cos
1-cos^2 = 4 - 8cos + 4cos^2
5cos^2 - 8cos + 3 = 0
cos = -0.8 ± 0.2
cos = -1, -0.6
Angle = 180º, 126.87º, 233.13º